# DocSLM: A Small Vision-Language Model for Long Multimodal Document Understanding

### v1

@@ 3 Method
Given a long multimodal document $\mathcal{D}=\{d^{1},d^{2},\dots,d^{N}\}$ with NN pages where $d^{n}$ is the $n^{th}$ page and a natural language query $S$, our goal is to predict a response $y$ that is consistent with the entire input. To achieve this under strict memory and context-length constraints, DocSLM introduces two key components: (1) a Hierarchical Multimodal Compression module that condenses visual, textual, and layout features into a compact token representation per page, and (2) a Streaming Abstention mechanism that enables reliable reasoning over arbitrarily long inputs. Overview of the method is in Fig. 4.
@@ 3.2 Streaming Abstention
Document Segmentation. For a sequence of length NN, attention memory scales linearly with sequence length, i.e., $\mathcal{O}(N)$ per layer when using optimized kernels such as FlashAttention [dao2022flashattention]. However, even linear growth becomes impractical on memory-constrained edge devices. To reduce the peak memory usage at any given time, we divide the document into $T$ smaller segments of equal length $N/T$ denoted as $\{s_{t}\}_{t=1}^{T}$.
@@ 5 Experiments
Comparison with Small Vision–Language Models. Within the 2–3B parameter range, DocSLM achieves the best trade-off between efficiency and accuracy (Tab.  2). Compared to Docopilot-2B [duan2025docopilot] (3,133 tokens, 35.9 ms), it runs faster (32.1 ms) while improving by +0.9 $p$ on MMLDoc and +26.3 $p$ on DUDE. Relative to InternVL2-2B [chen2024far], DocSLM gains +12.2 $p$ on MMLDoc and +47.6 $p$ on DUDE, highlighting its superior multimodal reasoning and text–layout alignment under tight token budgets.

### v2

@@ 3 Method
Given a long multimodal document $\mathcal{D}=\{d^{1},d^{2},\dots,d^{N}\}$ with $N$ pages where $d^{n}$ is the $n^{th}$ page and a natural language query $S$, our goal is to predict a response $y$ that is consistent with the entire input. To achieve this under strict memory and context-length constraints, DocSLM introduces two key components: (1) a Hierarchical Multimodal Compression module that condenses visual, textual, and layout features into a compact token representation per page, and (2) a Streaming Abstention mechanism that enables reliable reasoning over arbitrarily long inputs. Overview of the method is in Fig. 4.
@@ 3.2 Streaming Abstention
Document Segmentation. For a sequence of length $N$, attention memory scales linearly with sequence length, i.e., $\mathcal{O}(N)$ per layer when using optimized kernels such as FlashAttention [dao2022flashattention]. However, even linear growth becomes impractical on memory-constrained edge devices. To reduce the peak memory usage at any given time, we divide the document into $T$ smaller segments of equal length $N/T$ denoted as $\{s_{t}\}_{t=1}^{T}$.
@@ 5 Experiments
Comparison with Small Vision–Language Models. Within the 2–3B parameter range, DocSLM achieves the best trade-off between efficiency and accuracy (Tab.  2). Compared to Docopilot-2B [duan2025docopilot] (3,133 tokens, 35.9 ms), it runs faster (32.1 ms) while improving by +0.9 pp on MMLDoc and +26.3 pp on DUDE. Relative to InternVL2-2B [chen2024far], DocSLM gains +12.2 pp on MMLDoc and +47.6 pp on DUDE, highlighting its superior multimodal reasoning and text–layout alignment under tight token budgets.